A brick rests on a large piece of wood floating in a bucket of water. The brick slides off and sinks. Does the water level in the bucket go up, go down, or stay the same? Explain.
1
Expert's answer
2020-09-17T09:28:06-0400

Assume that the volume of the brick is V, its density is "\rho_b".

Initially, the brick was sitting on the wood with density and volume "\rho_w, V_w", and they were floating together. Assume that initially, the brick was not even touching water. "V_s" below is the volume of wood submerged under water. Since they are floating, the force of gravity is in balance with buoyancy force:


"g\rho_bV+g\rho_wV_w=g\rho V_s."

The volume of wood submerged below the water surface is


"V_s=\frac{\rho_bV}{\rho}+\frac{\rho_wV_w}{\rho}."

Initially, what makes water in bucket rise (compared to bucket with only water) was Vs only.

Finally, when the brick falls, the wood is submerged for


"V_{s2}=\frac{\rho_wV_w}{\rho}."

Finally, what makes water in bucket rise is


"V_{s2}+V=\frac{\rho_wV_w}{\rho}+V."

Sine density of brick is higher than density of water, in the equation for Vs above we have a coefficient "\rho_b/\rho>1" before brick volume.

In the last equation for Vs2, we see that the coefficient for V is 1. That is why the water level in the bucket will go down:

"V_s>V_{s2}+V."